OpenStudy (anonymous):

What value of c makes the expression y^2 -9y + c a perfect trunk mail square?

OpenStudy (gudden):

well.. break the middle term in the first step.. And divide it with 2 ( since (a+b)^2 = (a)^2 + 2ab + b^2) ( Where a is first term and B is second) so. to get ab we need to divide that term with 2 We get ab = (-9/2)y Since in the given equation a = y So, plugging in a as y in ab = (-9/2)y we get (y) (b) = (-9/2)y Now divide both sides of equation with y => b = -9/2 Hence value of c = b = -9/2.. Did you get it?